(a)Calculate all the four second-order partial derivatives of the following functions:
(i) f(x,y)=cos(x^2+y^2)
(ii) f(x,y)=sin(x/y)

(b) Find the range of the function f defined by f(x,y) 10–x^2–y^2 for all (x,y)
for which x^2+y^2 ≤9 Sketch two of its level curves.
(c)Check whether the following functions are differentiable at the point given
against them:
(i)f(x,y)=|x=1|at (1,0)
(ii) f(x,y) =y^3+ ysin2x +e^(x+y) at (1,–1)

(d)Find dw/dt
w = xy + x,z = cos t, y = sint , z=1 at t= 0 .
1
Expert's answer
2020-03-16T11:56:37-0400

"(a)(i)\nf(x,y)=cos(x ^\n2\n +y^ \n2\n )"

"f \n\u2032\n (x,y)_ \nx\n\u200b\t\n =\u22122xsin(x \n^2\n +y^ \n2\n )"

"f \n\u2032\u2032\n (x,y)_ \n{xx}\n\u200b\t\n =\u22122sin(x^ \n2\n +y ^\n2\n )\u22122x.2xcos(x^ \n2\n +y ^\n2\n )=\u22122sin(x^ \n2\n +y ^\n2\n )" "\u22124x^ \n2\n cos(x^ \n2\n +y ^\n2\n )"

"f \n\u2032\n (x,y)_ \ny\n\u200b\t\n =\u22122ysin(x ^\n2\n +y ^\n2\n )"

"f \n\u2032\u2032\n (x,y)_ \n{yy}\n\u200b\t\n =\u22122sin(x^ \n2\n +y ^\n2\n )\u22122y.2ycos(x^ \n2\n +y ^\n2\n )=\u22122sin(x^ \n2\n +y ^\n2\n )" "\u22124y^ \n2\n cos(x^ \n2\n +y ^\n2\n )"

"(ii) \ f(x,y)=sin(x/y)\\f'(x,y)_x=\dfrac{cos(x/y)}{y}\\f''(x,y)_{xx}=-\dfrac{sin(x/y)}{y^2} \\ f'(x,y)_y=-\frac{x\cos \left(\frac{x}{y}\right)}{y^2} \\f''(x,y)_{yy}=\mathrm{\:}x\frac{\frac{x\sin \left(\frac{x}{y}\right)}{y^2}y^2-2y\cos \left(\frac{x}{y}\right)}{\left(y^2\right)^2}=\frac{x\left(x\sin \left(\frac{x}{y}\right)-2y\cos \left(\frac{x}{y}\right)\right)}{y^4}"

"(b) f(x,y) =10–x^2–y^2\ \& \ \ \ \ ; x^2+y^2 ≤9\\ f(x,y)=10-(x^2+y^2) \\maximum \ value\ of x^2+y^2 \ is\ 9\ and \ minimum\ value\ is \ 0"

"f(x,y) \ ranges \ between \ 1 \ to\ 10"

"(c)f(x,y)=|x-1|\ at\ (0,1)\\"



From the graph we can see that there is sharp turn at x=1 which means the function is not differentiable at x=1

"(ii) y^{3}+y\sin2x+e^{x+y}=0"

In the above function no subpart of the function which is not differentiable at any point

sin , exponential and y^3 all are continuous and differentiable at all domain

Hence the above function is differentiable

"d) w = xy + x , x = cos t , y = sint"

substituting x and y in w

"w=cos t*sin t + cos t"

"w=(sin 2t)\/2 + cos t"

"dw\/dt= cos 2t - sin t"

"putting t=0\n\ndw\/dt= 1"